OpenStudy (zned6559):

Find the equation to the tangent line to the curve x^2 (2-x)-y^2 =0

OpenStudy (zned6559):

\[x ^{2}(2-x) - y ^{2}\]

OpenStudy (anonymous):

\[y=\sqrt{2x^2-x^3}\]

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=\frac{ 1\left( 4x-3x^2 \right) }{ 2\sqrt{2x^2-x^3} }\]

OpenStudy (anonymous):

find x,and then y eq. of tangent is y-?=dy/dx(x-?)